OpenStudy (anonymous):

a 300-n force acts on a 25-kg object. the acceleration of the object is?

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

f = m X a 300-n = 25 X a acceleration a = (300-n) / 25

OpenStudy (anonymous):

or do u mean a force of 300 newtons?

OpenStudy (anonymous):

if thats the case then a = 300/25 = 12 metres / ses/sec

OpenStudy (anonymous):

Using Newton's Second Law, (force) = (mass) x (acceleration). This is simply a plug-and-chug equation problem. (300 Newtons) = (25 kilograms) x (a) a = 12 meters/(second^2)

OpenStudy (anonymous):

why were u asking such a nice question. to check everybody's senses,i suppose.

OpenStudy (anonymous):

acc. to newton,s second law f=ma f=300N and mass=25kg so, 300 =25xa a=300/25=12m/secxsec

OpenStudy (anonymous):

F= m^a ie, 300n = 25 ^ a ie, 300/25 = a ie, 12=a therefore the acceleration is 12m/(sec)^2

OpenStudy (anonymous):

If we think about the object being in void, the mass of the object does not have relevance. Normally, in a gravitational field, an object falling from a certain height will have the formula F=m x g (where g is the gravitational acceleration=9.81 m/(s^2)). But, since there's no mention of such a thing, I will suppose that the object is in void. Being in void, its mass does not affect the object's acceleration. Thus, the problem does not have sense (since there is no gravitational field, and there is no more information).

OpenStudy (anonymous):

newtons second law f=ma.so a=f/m=300/25=12m/ss.very tough question.you got to thank me.